OpenStudy (anonymous):

find the slope of 7x=-4y

OpenStudy (anonymous):

Do you know the formula to use

OpenStudy (igreen):

You have to put it in the form of y = mx + b, where m is slope. 7x = -4y Add -4y to both sides: 7x + 4y = 0 Subtract 7x to both sides: 4y = -7x Divide 4 to both sides: y = -7/4x

OpenStudy (anonymous):

damn your smart

OpenStudy (igreen):

So the slope is -7/4.

OpenStudy (anonymous):

well, it easy but

OpenStudy (anonymous):

still

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @yancy234

OpenStudy (igreen):

Lol, thanks. @woooooooooowwwww

OpenStudy (anonymous):

np